OpenStudy (anonymous):

Two boxes of apples contain a total of 24 apples. If you halve the number of apples in the first box and add 4 apples to the second box, the total changes to 20 apples. How many apples are in each box initially?

OpenStudy (anonymous):

12 and 12 18 and 6 16 and 8 10 and 14

OpenStudy (anonymous):

Make two equations again x+y=24 \[\frac{ 1 }{ 2 }x+(y+4)=20\] subtract 4 on both sides and you get 1/2x+y=16 (x+y=24) - (1/2x+y=16)= \[\frac{ -1 }{ 2 }x=8\] Divide by -1/2 on both sides and x=16 x represents the first box and y represents the second box so 16+y=24 y=8 meaning there are 8 apples in the second box Answer is 16 and 8